An observer counts 360 vehicles per hour at a highway. Estimate the probability of having more than 4 vehicles arriving over a 25 second time interval.


1
Expert's answer
2021-05-31T04:11:02-0400

Solution:

"\lambda=" 360 vehicles per hour"=0.1" vehicles per second"=2.5" vehicles per 25-second

"X\sim Poi(\lambda)"

"P(X\ge4)=1-P(X<4)
\\=1-[P(X=0)+P(X=1)+P(X=2)+P(X=3)]"

"=1-[e^{-2.5}\dfrac{(2.5)^0}{0!}+e^{-2.5}\dfrac{(2.5)^1}{1!}+e^{-2.5}\dfrac{(2.5)^2}{2!}+e^{-2.5}\dfrac{(2.5)^3}{3!}]"

"=0.24242"